Carla runs 4 miles every day. she jobs from home to the school track, which is 3/4 mile away. she then runs laps around the 1/4 mile track. carla then jobs home. how many laps does she run at the school?
Answer by checkley77(12844)

4=3/4+3/4+1/4X
4=1.5+.25X
.25X=4-1.5
.25X=2.5
X=2.5/.25
X=10 LAPS.
PROOF:
4=3/4+3/4+1/4*10
4=1.5+10/4
4=1.5+2.5
4=4